# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 5.2 Factorization and two special subcases
${\cal C}(y(x))=\sum_{k=0}^{Q-1}{\rm e}^{-2xk}\prod_{j=0}^{k-1}\left(1+{\rm e}^{x-y(x)-\frac{\pi{\rm i}}{Q}(2j+1)}\right)\left(1+{\rm e}^{y(x)+\frac{\pi{\rm i}}{Q}(2j+1)}\right)\,,$ (5.202)
and $F_{0}(y(x))$ can be expressed explicitly by means of Eq. (C.13b). We observe that the reduced factorization formulae in Eq. (5.197), (5.200) for ${\mathsf{b}}^{2}=P$ and ${\mathsf{b}}^{2}=1/Q$ with $P,Q\in{\mathbb{Z}}_{>0}$ can alternatively be derived by substituting ${\rm i}{\mathsf{b}}k/P={\rm i}{\mathsf{b}}^{-1}k$ and ${\rm i}{\mathsf{b}}k/P={\rm i}{\mathsf{b}}k$ into the arguments of the Faddeev’s quantum dilogarithms in Eq. (5.186) and applying the periodicity properties in Eqs. (C.8c), (C.8d) and Eqs. (C.8a), (C.8b) to the resulting formulae for $F_{k}(y(x))$, $10\leq k\leq PQ-$ 1, respectively. In the general case of ${\mathsf{b}}^{2}=P/Q$, instead, we cannot resort to these quasi-periodicity properties because the functions $F_{k}(y(x))$ contain, in general, a fractional shift in the arguments of the Faddeev’s quantum dilogarithms.
@@ 6 Conclusions
In this paper, we have made various observations on the structure of the wave functions in CS theory with gauge group $\text{SL}(2,{\mathbb{C}})$ on the complement of a hyperbolic knot in the three-sphere. We have first conjectured explicit integrality properties for the resummed WKB expansion of the wave function, which can be tested directly using the decomposition into holomorphic/anti-holomorphic blocks. Our integrality conjecture guarantees that the singularities appearing in the holomorphic blocks at rational values of $\hbar$ get canceled against those occurring in the corresponding anti-holomorphic blocks. We have then developed various techniques to compute the wave function in the rational case effectively. In particular, we have analyzed the quantum $A$-polynomial at rational points by exploiting the underlying quasi-periodic structure, and we have solved the associated $q$-dfference equation in closed form in the general cases of order two and three. The calculation methods we have introduced have been applied to the examples of the figure-eight and the three-twist knots, and our conjectural statements have been verified. Finally, we have performed a direct evaluation of the Andersen–Kashaev state integral of the figure-eight knot at rational values, which gave us further insights into the properties of the exact wave functions. All computational approaches have yielded compatible results.
@@
Our investigation raises two main questions. The first one concerns the interpretation of the integrality structure of the resummed WKB expansion of the wave function in terms of an enumerative problem, or a BPS counting problem, in the dual supersymmetric theory dgg . The second one concerns the shape of the wave function obtained by directly evaluating the state integral. As highlighted in Section 5, this is enticingly similar to the functional profiles of the ground state wave functions computed in the context of the TS/ST correspondence mz-wv ; mz-wv2 . This observation raises the issue of understanding the AJ conjecture from a purely quantum mechanical point of view. The AJ conjecture can be explained, a priori, as a consequence of the quantization of the classical moduli space described by the $A$-polynomial, as pointed out in gukov and briefly recalled in Section 2. However, the precise details of this quantization procedure still need to be fully understood from a physical perspective. We know that the quantization is non-trivial and cannot be obtained by simply promoting the classical variables $X,P$ to Weyl operators. In addition, the classical phase space is not compact, which carries various problems concerning its appropriate quantization. From a quantum mechanical angle, in such a situation, we would generically expect to have at most metastable states with complex energies. We are, therefore, tempted to conjecture that the observed non-trivial quantization of the $A$-polynomial is required to obtain a normalizable wave function with precisely zero energy. We hope to address this question in the near future.

### v2

@@ 5.2 Factorization and two special subcases
${\cal C}(y(x))=\sum_{k=0}^{Q-1}{\rm e}^{-2xk}\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{x-y(x)-{{\pi{\rm i}\over Q}}(2j+1)}\bigr)\bigl(1+{\rm e}^{y(x)+{{\pi{\rm i}\over Q}}(2j+1)}\bigr),$ (5.33)
and $F_{0}(y(x))$ can be expressed explicitly by means of equation (C.20b). We observe that the reduced factorization formulae in equations (5.28) and (5.31) for ${\mathsf{b}}^{2}=P$ and ${\mathsf{b}}^{2}=1/Q$ with $P,Q\in{\mathbb{Z}}_{>0}$} can alternatively be derived by substituting ${\rm i}{\mathsf{b}}k/P={\rm i}{\mathsf{b}}^{-1}k$ and ${\rm i}{\mathsf{b}}k/P={\rm i}{\mathsf{b}}k$ into the arguments of the Faddeev’s quantum dilogarithms in equation (5.17) and applying the periodicity properties in equations (C.15c) and (C.15d) and equations (C.15a) and (C.15b) to the resulting formulae for $F_{k}(y(x))$, $0\leq k\leq PQ-1$, respectively. In the general case of ${\mathsf{b}}^{2}=P/Q$, instead, we cannot resort to these quasi-periodicity properties because the functions $F_{k}(y(x))$ contain, in general, a fractional shift in the arguments of the Faddeev’s quantum dilogarithms.
@@ 6 Conclusions
In this paper, we have made various observations on the structure of the wave functions in CS theory with gauge group $\mathrm{SL}(2,{\mathbb{C}})$ on the complement of a hyperbolic knot in the three-sphere. We have first conjectured explicit integrality properties for the resummed WKB expansion of the wave function, which can be tested directly using the decomposition into holomorphic/antiholomorphic blocks. Our integrality conjecture guarantees that the singularities appearing in the holomorphic blocks at rational values of $\hbar$ get canceled against those occurring in the corresponding antiholomorphic blocks. We have then developed various techniques to compute the wave function in the rational case effectively. In particular, we have analyzed the quantum $A$-polynomial at rational points by exploiting the underlying quasi-periodic structure and solved the associated $q$-difference equation in closed form in the general cases of orders two and three. The calculation methods we have introduced have been applied to the examples of the figure-eight and three-twist knots, and our conjectural statements have been verified. Finally, we have performed a direct evaluation of the Andersen–Kashaev state integral of the figure-eight knot at rational values, which gave us further insights into the properties of the exact wave functions. All computational approaches yielded compatible results.
@@
Our investigation raises two main questions. The first one concerns the interpretation of the integrality structure of the resummed WKB expansion of the wave function in terms of an enumerative problem, or a BPS counting problem, in the dual supersymmetric theory [10], either directly or potentially exploiting results from the resurgent analysis of the perturbative wave functions performed in [17]. The second one concerns the shape of the wave function obtained by directly evaluating the state integral. As highlighted in Section 5, this is enticingly similar to the functional profiles of the ground state wave functions computed in the context of the TS/ST correspondence [38, 39]. This observation raises the issue of understanding the AJ conjecture of [16] from a purely quantum-mechanical point of view. The AJ conjecture can be explained, a priori, as a consequence of the quantization of the classical moduli space described by the $A$-polynomial, as pointed out in [22] and briefly recalled in Section 2. However, the precise details of this quantization procedure still need to be fully understood from a physical perspective. We know that the quantization is non-trivial and cannot be obtained by simply promoting the classical variables $X$, $P$ to Weyl operators. In addition, the classical phase space is not compact, which carries various problems concerning its appropriate quantization. From a quantum-mechanical angle, in such a situation, we would generically expect to have, at most, metastable states with complex energies. We are, therefore, tempted to conjecture that the observed non-trivial quantization of the $A$-polynomial is required to obtain a normalizable wave function with precisely zero energy. We hope to address this question in the near future.
Appendix A Computing the $\bm{\epsilon}$-expansions
For completeness, we include here the details of the computations performed in Section 3.2. We observe that, as a consequence of equations (3.3) and (3.7), the coefficient functions $\hat{a}_{\pm n}(q)$ satisfy the relation
$\hat{a}_{\pm n}(q)q^{-\sigma_{\pm n}}=\sum_{k|n}kD_{\pm k}\bigl(q^{n\over k}\bigr)q^{-{n\over k}\sigma_{\pm k}}\sum_{j=0}^{k-1}q^{{n\over k}j},\qquad n\in{\mathbb{Z}}_{>0},$ (A.1)
and we write the $\mathbb{Z}$}-polynomials $D_{\pm k}(q)$ explicitly as
$D_{\pm k}(q)=\sum_{\ell=0}^{\deg{D_{\pm k}}}D_{\pm k,\ell}q^{\ell},\qquad k\in{\mathbb{Z}}_{>0},$ (A.2)
$q={\rm e}^{2\pi{\rm i}{P\over Q}},\qquad P,Q\in{\mathbb{Z}}_{>0}\ {\rm coprime},$ (A.3)
while we introduce
$q_{\epsilon}={\rm e}^{{\rm i}\hbar_{\epsilon}},\qquad\hbar_{\epsilon}=2\pi{P\over Q}+\epsilon,\qquad 0<\epsilon\ll 1,$ (A.4)
$D_{\pm k}\bigl(q_{\epsilon}^{sQ\over k}\bigr)=\left(1+{\rm i}\epsilon q{\partial\over\partial q}\right)D_{\pm k}\bigl(q^{sQ\over k}\bigr)+{\cal O}\bigl(\epsilon^{2}\bigr),$ (A.5a)
$\sum_{j=0}^{k-1}q_{\epsilon}^{{sQ\over k}(j-\sigma_{\pm k})}=\left(1+{\rm i}\epsilon q{\partial\over\partial q}\right)\sum_{j=0}^{k-1}q^{{sQ\over k}(j-\sigma_{\pm k})}+{\cal O}\bigl(\epsilon^{2}\bigr),$ (A.5b)
and therefore also
$\hat{a}_{\pm sQ}(q_{\epsilon})q_{\epsilon}^{-\sigma_{\pm sQ}}=\left(1+{\rm i}\epsilon q{\partial\over\partial q}\right)\hat{a}_{\pm sQ}(q)q^{-\sigma_{\pm sQ}}+{\cal O}\bigl(\epsilon^{2}\bigr).$ (A.6)
We now observe that
${{1\over q_{\epsilon}^{sQ}-1}}=-{{{\rm i}\over sQ\epsilon}}-{{1\over 2}}+{\cal O}(\epsilon),$ (A.7)
and, combining equations (A.6) and (A.7), we finally obtain the NLO $\epsilon$-expansion in equation (3.22a). On the dual side, we consider
$q_{D}=q^{Q^{2}\over P^{2}}={\rm e}^{2\pi{\rm i}{Q\over P}},\qquad q_{\epsilon,D}={\rm e}^{4\pi^{2}{\rm i}\over\hbar_{\epsilon}},$ (A.8)
and take again the limit $\epsilon\rightarrow 0$. It follows that
$q_{\epsilon,D}=q_{D}{\rm e}^{-{\rm i}\epsilon{Q^{2}\over P^{2}}+{\cal O}(\epsilon^{2})}.$ (A.9)
$D_{\pm k}\Big(q_{\epsilon,D}^{sP\over k}\Big)=\left(1-{\rm i}\epsilon{{Q^{2}\over P^{2}}}q_{D}{\partial\over\partial q_{D}}\right)D_{\pm k}\Big(q_{D}^{sP\over k}\Big)+{\cal O}\bigl(\epsilon^{2}\bigr),$ (A.10a)
$\sum_{j=0}^{k-1}q_{\epsilon,D}^{{sP\over k}(j-\sigma_{\pm k})}=\left(1-{\rm i}\epsilon{{Q^{2}\over P^{2}}}q_{D}{\partial\over\partial q_{D}}\right)\sum_{j=0}^{k-1}q_{D}^{{sP\over k}(j-\sigma_{\pm k})}+{\cal O}\bigl(\epsilon^{2}\bigr),$ (A.10b)
and therefore also
$\hat{a}_{\pm sP}(q_{\epsilon,D})q_{\epsilon,D}^{-\sigma_{\pm sP}}=\left(1-{\rm i}\epsilon{{Q^{2}\over P^{2}}}q_{D}{\partial\over\partial q_{D}}\right)\hat{a}_{\pm sP}(q_{D})q_{D}^{-\sigma_{\pm sP}}+{\cal O}\bigl(\epsilon^{2}\bigr).$ (A.11)
Finally, we note that
${{1\over q_{\epsilon,D}^{sP}-1}}={{{\rm i}P\over sQ^{2}\epsilon}}-{{1\over 2}}+{{{\rm i}\over 2\pi sQ}}+{\cal O}(\epsilon),$ (A.12)
and, combining equations (A.11) and (A.12), we obtain the NLO $\epsilon$-expansion in equation (3.22b).
Appendix B An alternative way of obtaining the integer sequences
We present an alternative, simple way of computing the integer sequences $\{c_{\pm s}\}$, $s\in{\mathbb{Z}}_{>0}$, defined in equations (3.27) and (3.28), using the results on the exact wave function at rational values obtained in Section 3.3. Let us denote by $\phi_{(\pm)}^{\rm WKB}(X;q)$ the contributions to the holomorphic component $\phi^{\rm WKB}(X;q)$ from the positive and negative powers of $X$, respectively, so that equation (3.2) assumes the form
and analogously for the antiholomorphic component $\phi^{\rm WKB}(X_{D};q_{D})$. We fix $\hbar$ as in equation (3.15) and consider the formula for $\phi^{\rm WKB}(X;q)+\phi^{\rm WKB}(X_{D};q_{D})$ in equation (3.44). Setting the values $P=Q=1$, the $X$-series becomes
${{{\rm i}\over 2\pi}}\sum_{s\geq 1}{{X^{s}\over s^{2}}}\left[(2\pi{\rm i}s+1-s\log X+4\pi{\rm i}\sigma_{s})\hat{a}_{s}(1)-4\pi{\rm i}{{\partial\over\partial q}}\hat{a}_{s}(q)_{q=1}\right].$ (B.2)
Let us now go back to the sum in equation (3.16) and apply the change of variable $\hbar\rightarrow$-$\hbar$, which transforms $q$, $q_{D}$ into $q^{-1}$, $q_{D}^{-1}$, while keeping $X$, $X_{D}$ fixed. After performing the $\epsilon$-expansion as in Section 3.2 and then taking $P=Q=1$ in the sum $\phi_{(+)}^{\rm WKB}\bigl(X;q^{-1}\bigr)+\phi_{(+)}^{\rm WKB}\bigl(X_{D};q_{D}^{-1}\bigr)$, we find the very close but not identical quantity
${{{\rm i}\over 2\pi}}\sum_{s\geq 1}{{X^{s}\over s^{2}}}\left[(2\pi{\rm i}s-1+s\log X+4\pi{\rm i}\sigma_{s})\hat{a}_{s}(1)-4\pi{\rm i}{{\partial\over\partial q}}\hat{a}_{s}(q)_{q=1}\right].$ (B.3)
By taking the difference of equations (B.2) and (B.3), we arrive at
${{{\rm i}\over\pi}}\sum_{s\geq 1}{{X^{s}\over s^{2}}}(1-s\log X)\hat{a}_{s}(1).$ (B.4)
Therefore, if we compute the function
$\big[\bigl(\phi_{(+)}^{\rm WKB}(X;q)+\phi_{(+)}^{\rm WKB}(X_{D};q_{D})\bigr)-\bigl(\phi_{(+)}^{\rm WKB}\bigl(X;q^{-1}\bigr)+\phi_{(+)}^{\rm WKB}\bigl(X_{D};q_{D}^{-1}\bigr)\bigr)\big]_{P=Q=1},$ (B.5)
its expansion in powers of $X$ is precisely the series in equation (B.4). From the coefficients of this expansion, we can easily extract the desired integers $c_{s}=\hat{a}_{s}(1)$ for $s\in{\mathbb{Z}}_{>0}$. Let us now consider the $X^{-1}$-series in equation (3.44). Again, setting the values $P=Q=1$, we find
${{{\rm i}\over 2\pi}}\sum_{s\geq 1}{{X^{-s}\over s^{2}}}\left[(2\pi{\rm i}s+1+s\log X)\hat{a}_{-s}(1)-4\pi{\rm i}{{\partial\over\partial q}}\hat{a}_{-s}(q)_{q=1}\right].$ (B.6)
As before, we go back to the sum in equation (3.16) and apply the change of variable $\hbar\rightarrow$-$\hbar$, which transforms $q$, $q_{D}$ into $q^{-1}$, $q_{D}^{-1}$, while keeping $X$, $X_{D}$ fixed. After performing the $\epsilon$-expansion as in Section 3.2 and then taking $P=Q=1$ in the sum
$\phi_{(-)}^{\rm WKB}\bigl(X;q^{-1}\bigr)+\phi_{(-)}^{\rm WKB}\bigl(X_{D};q_{D}^{-1}\bigr),$
we find the slightly different quantity
${{{\rm i}\over 2\pi}}\sum_{s\geq 1}{{X^{-s}\over s^{2}}}\left[(2\pi{\rm i}s-1-s\log X)\hat{a}_{-s}(1)-4\pi{\rm i}{{\partial\over\partial q}}\hat{a}_{-s}(q)_{q=1}\right].$ (B.7)
By taking the difference of equations (B.6) and (B.7), we arrive at
${{{\rm i}\over\pi}}\sum_{s\geq 1}{{X^{-s}\over s^{2}}}(1+s\log X)\hat{a}_{-s}(1).$ (B.8)
Accordingly, if we compute the function
$\big[\bigl(\phi_{(-)}^{\rm WKB}(X;q)+\phi_{(-)}^{\rm WKB}(X_{D};q_{D})\bigr)-\bigl(\phi_{(-)}^{\rm WKB}\bigl(X;q^{-1}\bigr)+\phi_{(-)}^{\rm WKB}\bigl(X_{D};q_{D}^{-1}\bigr)\bigr)\big]_{P=Q=1},$ (B.9)
its expansion in powers of $X^{-1}$ is precisely the series in equation (B.8). From the coefficients of this expansion, we can straightforwardly obtain the desired integers $c_{-s}=\hat{a}_{-s}(1)$ for $s\in{\mathbb{Z}}_{>0}$.
The quantum dilogarithm is the function of two variables defined by the series [14, 32]
$(zq^{\alpha};q)_{\infty}=\prod_{i=0}^{\infty}\bigl(1-zq^{\alpha+i}\bigr),\qquad\alpha\in{\mathbb{R}},$ (C.1)
which is analytic in $z,q\in{\mathbb{C}}$ with $|q|<1$ and admits asymptotic expansions around $q$ a root of unity. The $q$-Pochhammer symbols, also known as $q$-shifted factorials, are defined by
$(z;q)_{n}=\prod_{i=0}^{n-1}\bigl(1-zq^{i}\bigr),\qquad n\in{\mathbb{Z}}_{>0},$ (C.2a)
$(z;q)_{-n}={{1\over(zq^{-n};q)_{n}}}={{(-z)^{-n}q^{n(n+1)/2}\over\bigl(z^{-1}q;q\bigr)_{n}}}={{1\over\bigl(zq^{-1};q^{-1}\bigr)_{n}}},\qquad n\in{\mathbb{Z}}_{>0},$ (C.2b)
with $(z;q)_{0}=1$. Equivalently, we can write
$(z;q)_{n}={{(z;q)_{\infty}\over(zq^{n};q)_{\infty}}},\qquad n\in{\mathbb{Z}}.$ (C.3)
Moreover, the quantum dilogarithm satisfies the $q$-binomial theorem, which can be expressed as
${{(xz;q)_{\infty}\over(z;q)_{\infty}}}=\sum_{n=0}^{\infty}{{(x;q)_{n}\over(q;q)_{n}}}z^{n},$ (C.4)
where $x,z,q\in{\mathbb{C}}$ and $|q|<1$, and implies the following $q$-series identities:
$(qz;q)_{\infty}=\sum_{n=0}^{\infty}{{(-1)^{n}q^{n(n+1)/2}\over(q;q)_{n}}}z^{n}={{1\over\bigl(z;q^{-1}\bigr)_{\infty}}},$ (C.5a)
${{1\over(z;q)_{\infty}}}=\sum_{n=0}^{\infty}{{z^{n}\over(q;q)_{n}}}=\bigl(q^{-1}z;q^{-1}\bigr)_{\infty}.$ (C.5b)
The Faddeev’s quantum dilogarithm $\Phi_{{\mathsf{b}}}(z)$ is defined in the strip $|\operatorname{Im}(z)|<|\operatorname{Im}(c_{{\mathsf{b}}})|$, where
$c_{{\mathsf{b}}}={\rm i}\bigl({\mathsf{b}}+{\mathsf{b}}^{-1}\bigr)/2,$ (C.6)
by the integral representation [13, 14]
$\Phi_{{\mathsf{b}}}(z)=\exp\left(\int_{{\mathbb{R}}+{\rm i}\epsilon}{{{\rm e}^{-2{\rm i}zy}\over 4\sinh(y{\mathsf{b}})\sinh\bigl(y{\mathsf{b}}^{-1}\bigr)}}{{{\rm d}y\over y}}\right),$ (C.7)
which implies the symmetry properties
$\Phi_{{\mathsf{b}}}(z)=\Phi_{-{\mathsf{b}}}(z)=\Phi_{{\mathsf{b}}^{-1}}(z).$ (C.8)
When $\operatorname{Im}\big({\mathsf{b}}^{2}\big)>0$, the formula in equation (C.7) is equivalent to
$\Phi_{{\mathsf{b}}}(z)={{\big({\rm e}^{2\pi{\mathsf{b}}(z+c_{{\mathsf{b}}})};q\big)_{\infty}\over\big({\rm e}^{2\pi{\mathsf{b}}^{-1}(z-c_{{\mathsf{b}}})};\tilde{q}\big)_{\infty}}}=\prod_{n=0}^{\infty}{{1-{\rm e}^{2\pi{\mathsf{b}}(z+c_{{\mathsf{b}}})}q^{n}\over 1-{\rm e}^{2\pi{\mathsf{b}}^{-1}(z-c_{{\mathsf{b}}})}\tilde{q}^{n}}},$ (C.9)
where
$q={\rm e}^{2\pi\mathrm{i}{\mathsf{b}}^{2}},\qquad\tilde{q}={\rm e}^{-2\pi\mathrm{i}{\mathsf{b}}^{-2}}.$ (C.10)
Note that the function in equation (C.9) can be extended to the region $\operatorname{Im}\bigl({\mathsf{b}}^{2}\bigr)<0$ by means of equation (C.8) and further admits an analytic continuation to all values of $\mathsf{b}$} such that ${\mathsf{b}}^{2}\notin{\mathbb{R}}_{\leq 0}$. Moreover, $\Phi_{{\mathsf{b}}}(z)$ can be extended to the whole complex $z$-plane as a meromorphic function with an essential singularity at infinity, poles at the points
$z=c_{{\mathsf{b}}}+{\rm i}m{\mathsf{b}}+{\rm i}n{\mathsf{b}}^{-1},$ (C.11)
and zeros at the points
$z=-c_{{\mathsf{b}}}-{\rm i}m{\mathsf{b}}-{\rm i}n{\mathsf{b}}^{-1},$ (C.12)
for $m,n\in{\mathbb{N}}$. It satisfies the inversion formula
$\Phi_{{\mathsf{b}}}(z)\Phi_{{\mathsf{b}}}(-z)={\rm e}^{\pi{\rm i}z^{2}}\Phi_{{\mathsf{b}}}(0)^{2},\qquad\Phi_{{\mathsf{b}}}(0)=\left({{q\over\tilde{q}}}\right)^{1/48}={\rm e}^{\pi{\rm i}({\mathsf{b}}^{2}+{\mathsf{b}}^{-2})/24},$ (C.13)
the complex conjugation formula
$\overline{\Phi_{{\mathsf{b}}}(z)}={{1\over\Phi_{\overline{{\mathsf{b}}}}(\overline{z})}},$ (C.14)
and is a quasi-periodic function. Precisely, as a direct consequence of its definition in equation (C.9), we find the relations
$\Phi_{{\mathsf{b}}}(z+{\rm i}{\mathsf{b}}k)=\Phi_{{\mathsf{b}}}(z)\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{2\pi{\mathsf{b}}z+\pi{\rm i}{\mathsf{b}}^{2}(2j+1)}\bigr)^{-1},$ (C.15a)
$\Phi_{{\mathsf{b}}}(z-{\rm i}{\mathsf{b}}k)=\Phi_{{\mathsf{b}}}(z)\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{2\pi{\mathsf{b}}z-\pi{\rm i}{\mathsf{b}}^{2}(2j+1)}\bigr),$ (C.15b)
$\Phi_{{\mathsf{b}}}\bigl(z+{\rm i}{\mathsf{b}}^{-1}k\bigr)=\Phi_{{\mathsf{b}}}(z)\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{2\pi{\mathsf{b}}^{-1}z+\pi{\rm i}{\mathsf{b}}^{-2}(2j+1)}\bigr)^{-1},$ (C.15c)
$\Phi_{{\mathsf{b}}}\bigl(z-{\rm i}{\mathsf{b}}^{-1}k\bigr)=\Phi_{{\mathsf{b}}}(z)\prod_{j=0}^{k-1}\bigl(1+{\rm e}^{2\pi{\mathsf{b}}^{-1}z-\pi{\rm i}{\mathsf{b}}^{-2}(2j+1)}\bigr),$ (C.15d)
${\mathsf{b}}^{2}={{P\over Q}},\qquad P,Q\in{\mathbb{Z}}_{>0}\ {\rm coprime},$ (C.16)
then the periodicity formulae above give in particular
$\Phi_{{\mathsf{b}}}(z+{\rm i}{\mathsf{b}}Q)$ $=\Phi_{{\mathsf{b}}}(z)\bigl(1+{\rm e}^{2\pi{\mathsf{b}}Qz+\pi{\rm i}PQ}\bigr)^{-1},$ (C.17a)
$\Phi_{{\mathsf{b}}}(z-{\rm i}{\mathsf{b}}Q)$ $=\Phi_{{\mathsf{b}}}(z)\bigl(1+{\rm e}^{2\pi{\mathsf{b}}Qz-\pi{\rm i}PQ}\bigr)^{.}$ (C.17b)
Finally, in the above rational case of equation (C.16), it was shown in [18] that Faddeev’s quantum dilogarithm can be expressed as
$\Phi_{{\mathsf{b}}}\left({{z\over 2\pi{\mathsf{b}}}}\right)={}$ $\exp\left[{{{\rm i}\over 2\pi PQ}}\right.\operatorname{Li}_{2}\bigl({\rm e}^{w(z)}\bigr)+\left(1+{{{\rm i}w(z)\over 2\pi PQ}}\right)\log\bigl(1-{\rm e}^{w(z)}\bigr)$
$\left.-\log D_{Q}\bigl({\rm e}^{{{w(z)\over Q}}};{\rm e}^{2\pi{\rm i}{{P\over Q}}}\bigr)-\log D_{P}\bigl({\rm e}^{{{w(z)\over P}}};{\rm e}^{2\pi{\rm i}{{Q\over P}}}\bigr)\right],$ (C.18)
where $w(z)=Qz+\pi{\rm i}(P+Q)$, $\operatorname{Li}_{2}(z)$ is the standard dilogarithm, and
$D_{N}(z;q)=\prod_{j=1}^{N-1}\bigl(1-zq^{j}\bigr)^{j/N},\qquad N\in{\mathbb{Z}}_{>0},$ (C.19)
is the $N$-th cyclic quantum dilogarithm with $D_{1}(z;q)=1$. In the special subcases of ${\mathsf{b}}^{2}=P$ positive integer and ${\mathsf{b}}^{2}=1/Q$ positive unit fraction, the formula in equation (C.18) simplifies into
$\Phi_{{\mathsf{b}}}\left({{z\over 2\pi{\mathsf{b}}}}\right)={}$ $\exp\left[{{{\rm i}\over 2\pi P}}\operatorname{Li}_{2}\bigl(-{\rm e}^{z+\pi{\rm i}P}\bigr)+\left({{P-1\over 2P}}+{{{\rm i}z\over 2\pi P}}\right)\log\bigl(1+{\rm e}^{z+\pi{\rm i}P}\bigr)\right$.
$\left.-\log D_{P}\bigl(-{\rm e}^{{{z\over P}}+{{\pi{\rm i}\over P}}};{\rm e}^{{{2\pi{\rm i}\over P}}}\bigr)\right],$ (C.20a)
$\Phi_{{\mathsf{b}}}\left({{z\over 2\pi{\mathsf{b}}}}\right)={}$ $\exp\left[{{{\rm i}\over 2\pi Q}}\operatorname{Li}_{2}\bigl(-{\rm e}^{Qz+\pi{\rm i}Q}\bigr)+\left({{Q-1\over 2Q}}+{{{\rm i}z\over 2\pi}}\right)\log\bigl(1+{\rm e}^{Qz+\pi{\rm i}Q}\bigr)\right$.
$\left.-\log D_{Q}\bigl(-{\rm e}^{z+{{\pi{\rm i}\over Q}}};{\rm e}^{{{2\pi{\rm i}\over Q}}}\bigr)\right],$ (C.20b)
respectively.